# Interpretable by Design: Wrapper Boxes Combine Neural Performance with Faithful Explanations and Example-based Model Auditing

### v2

@@ Abstract
Can we preserve the accuracy of neural models while also providing faithful explanations of model decisions with respect to training data? We propose “wrapper boxes”: training a neural model as usual and then using its learned feature representation in a classic, interpretable model to perform prediction. Across three large pre-trained language models, two datasets of varying scales, three classic models, and four evaluation metrics, we first show that predictive performance of wrapper boxes is largely comparable to the original neural models. Because our classic models are interpretable by design, each model decision is determined by a known set of training examples that can be directly shown to users. Wrapper boxes thus preserve the predictive performance of neural models while also faithfully attributing model decisions to training data. Among other use cases, such attribution enables model decisions to be contested on the basis of the responsible training examples. In comparison to prior work, we show our approach achieves both higher coverage and correctness in identifying which training data to remove in order to change a model decision. Our source code will be shared at publication for reproducibility.
Interpretable by Design: Wrapper Boxes Combine Neural Performance with Faithful Attribution of Model Decisions to Training Data
@@
Figure 1: Three wrapper boxes illustrated for toxic language detection (Hartvigsen et al., 2022). Red and blue dots denote harmful vs. benign speech. Smaller dots represent examples, while larger dots represent clusters (e.g., DT leaf nodes). A neural model’s penultimate layer provides the feature representation for the wrapper boxes. Our results show that wrapper boxes achieve comparable performance to the underlying neural models while also providing intuitive, example-based explanations via interpretable by design wrapper models (described in Section 4).
@@ 1 Introduction
Opaque predictive models are difficult to trust and reason about, with users citing the need for automated decisions to be transparent and interpretable (Langer et al., 2021; Shin, 2021). In critical sectors like law, health, and finance, interpretability may be essential to prevent catastrophic failures (Ahmad et al., 2018; Rudin, 2019; Bhatt et al., 2020). Furthermore, interpretability may be required for regulatory compliance (Kaminski, 2019). However, popular pre-trained language models (Devlin et al., 2019; Lewis et al., 2020; Floridi and Chiriatti, 2020; Chung et al., 2022) are inscrutable, making it very difficult to explain model decisions (Adadi and Berrada, 2018; Barredo Arrieta et al., 2020).
In contrast, classic “white box” methods such as $k$-nearest neighbor ($k$NN) and decision tree (DT) are inherently interpretable, or interpretable by design Rudin (2019): each model decision is determined by a known set of training examples that can be directly shown to users. However, classic models tend to underperform today’s neural models.
@@
In addition, research on interpretable NLP has largely focused on attribute-style explanations, with far less work on example-based explanations (Keane and Kenny, 2019). Because people naturally reason by analogy Sørmo et al. (2005); Schank et al. (2014); Kolodner (2014), explaining predictions in relation to specific training data is intuitively appealing. Example-based explanations also connect to traditional AI research on case-based reasoning Aamodt and Plaza (1994), relating new problem instances to similar past ones, a problem-solving strategy that people naturally use in decision-making (Newel and Simon, 1972). Rudin et al. (2022) thus argues for developing modern case-based methods as one of the grand challenges in interpretable machine learning today.
In this work, we propose wrapper boxes that combine the accuracy of modern neural models with the faithful, example-based explanations of classic, interpretable by design models. Specifically, we “wrap” a given neural model with one or more white box classic models. Building on the tradition of fitting fully connected layers on neural representations Krizhevsky et al. (2012); Simonyan and Zisserman (2015); He et al. (2016); Devlin et al. (2019), we fit white classic models on extracted neural representations for inference. The reasoning process of a wrapper box can then be faithfully explained by showing the specific training examples that led to each wrapper box prediction (Figure 1).
Note that wrapper boxes do not attempt to approximate the underlying neural model, i.e., wrapper box vs. neural model predictions often differ. Our claim is that wrapper box predictions can be faithfully explained and that the predictive performance of wrapper boxes is largely comparable to the underlying neural models (see below).
color=blue!20]Sam: Let’s add a sentence here about faithfulness
@@
We evaluate another such use case: enabling model decisions to be contested based on the training data responsible for those decisions. More specifically, we consider the task of identifying which training data would need to be removed in order to change a model decision Yang et al. (2023). color=orange!40]Jessy: minor edits here. This task offers a form of algorithmic recourse (Karimi et al., 2022), which emphasizes providing actionable explanations to users unfavorably treated by automated systems. Users are provided a foundation for contesting model decisions by attributing model decisions to specific training data. Compared to Yang et al. (2023), we show higher coverage and correctness in identifying which training data to remove while also generalizing beyond the simple linear models considered in their work.
@@ 2 Related Work
Case Inference Examples Explanation Examples Performance Faithfulness Simplicity
@@ III Subset Subset $\circ$ $\checkmark$ $\checkmark$
Table 2: Interpretable by Design models permit tradeoffs between key outcome variables – predictive performance, explanation faithfulness (or fidelity), and explanation simplicity – based on which training examples are used to make a prediction vs. to explain that prediction. Note that for any given input, different models will naturally vary in which training examples they consult as relevant to perform (and faithully explain) inference for that input.
@@ 3 Example-based Explanation Tradeoffs
We focus on interpretable by design predictive models that tie inference directly to specific training examples, enabling each prediction to be faithfully explained via those same training examples that determined the model’s prediction. Appendix D further discusses user perceptions of machine-retrieved examples. color=blue!20]Sam: appendix mention here
To better elucidate the design space for working with such models, this section illustrates possible tradeoffs between three key variables of interest: predictive performance, explanation faithfulness, and explanation simplicity. Section 4 discusses how such tradeoffs can be operationalized in practice for our specific wrapper box models.
@@ 4 Wrapper Boxes
Wrapper Boxes “wrap” a given neural model with one or more white box classic models by fitting each white classic model on extracted neural representations for inference. Note that wrapper boxes do not attempt to faithfully approximate the underlying neural model. While both approaches leverage learned linearly separable neural representations, the underlying decision-making process differs. Hence, derived example-based explanations faithfully explain the inference procedure of the wrapper boxes, not the original neural model.
Post hoc methods evaluate fidelity with respect to the neural model they seek to explain DeYoung et al. (2020); Jacovi and Goldberg (2020) since explanations can diverge from actual model behavior. In contrast, wrapper boxes are interpretable by design, where derived example-based explanations by construction must have been consulted during inference. Loss in fidelity can only occur intentionally if fewer training examples are shown in the explanation to reduce information overload.
@@ 4.1 Learning Feature Representations
As shown in Figure 1, we start with a fine-tuned neural model that acts as a task-specific encoder to learn high-quality embeddings for the input text. color=blue!20]Sam: Small nod to prior approaches here Whereas traditional neural models often fit linear classifiers on learned representations, we extract these representations for use by various classic, white box classifiers. This substitution thus enables prediction supported by faithful, example-based explanations and is agnostic to the neural architecture, training procedure, and data used.
@@ 4.2 Wrapper Box Models
We consider three interpretable by design models in which inference is directly linked to training examples. This means that, by design, model predictions can be faithfully and intuitively attributed to specific relevant training examples.
@@ 7 Conclusion
Wrapper boxes provide faithful, example-based explanations for classic model decisions, attributing decisions to specific training data. They are quite general and agnostic to the underlying neural architecture, training procedure, and input data. After training a neural model as usual, the learned feature representation is input to an interpretable by design Rudin (2019) classic model to perform prediction. Because classic models tie inference directly to specific training data, each prediction can be faithfully attributed to the training examples responsible.
Our first evaluation showed that wrapper boxes deliver predictive performance largely comparable to baseline transformers, as seen across three large pre-trained language models, two datasets of varying scale, three classic models, and four metrics.
We then discussed how such attribution enables model decisions to be contested based on the training data responsible for those decisions. In comparison to prior work Yang et al. (2023), we show our approach achieves both higher coverage and correctness in identifying which training data to remove in order to change a model decision.
@@ Acknowledgments
This research was supported in part by Cisco and by Good Systems  [^1] http://goodsystems.utexas.edu/, a UT Austin Grand Challenge to develop responsible AI technologies. Our opinions reflect our own views only.

### v3

# Wrapper Boxes: Faithful Attribution of Model Predictions to Training Data

@@ Abstract
Can we preserve the accuracy of neural models while also providing faithful explanations of model decisions to training data? We propose a “wrapper box” pipeline: training a neural model as usual and then using its learned feature representation in classic, interpretable models to perform prediction. Across seven language models of varying sizes, including four large language models (LLMs), two datasets at different scales, three classic models, and four evaluation metrics, we first show that predictive performance of wrapper classic models is largely comparable to the original neural models.
Because classic models are transparent, each model decision is determined by a known set of training examples that can be directly shown to users. Our pipeline thus preserves the predictive performance of neural language models while faithfully attributing classic model decisions to training data. Among other use cases, such attribution enables model decisions to be contested based on responsible training instances. Compared to prior work, our approach achieves higher coverage and correctness in identifying which training data to remove to change a model decision. To reproduce findings, our source code is online at: https://github.com/SamSoup/WrapperBox.
Wrapper Boxes: Faithful Attribution of Model Predictions to Training Data
@@
Figure 1: Three wrapper boxes illustrated for toxic language detection (Hartvigsen et al., 2022). Red and blue dots denote harmful vs. benign speech. Smaller dots represent examples, while larger dots represent clusters (e.g., DT leaf nodes). A neural model’s penultimate layer provides the feature representation for the white wrapper boxes. Our results show that classic models can achieve comparable performance to the underlying neural models while also providing intuitive, example-based explanations (described in Section 4).
@@ 1 Introduction
Opaque predictive models are challenging to trust and reason about, prompting calls for greater transparency and interpretability in automated decisions (Langer et al., 2021; Shin, 2021). In critical sectors like law, health, and finance, interpretability may be essential to prevent catastrophic failures (Ahmad et al., 2018; Rudin, 2019; Bhatt et al., 2020). Furthermore, interpretability may be required for regulatory compliance (Kaminski, 2019). However, popular pre-trained language models (Devlin et al., 2019; Lewis et al., 2020; Floridi and Chiriatti, 2020; Chung et al., 2022) are inscrutable, making it difficult to explain model decisions (Adadi and Berrada, 2018; Barredo Arrieta et al., 2020).
In contrast, classic “white box” methods such as $k$-nearest neighbor ($k$NN) and decision tree (DT) are inherently interpretable Rudin (2019): each model decision is determined by a known set of training examples that can be directly shown to users. Nevertheless, classic models tend to underperform today’s neural models.
@@
In addition, research on interpretable NLP has largely focused on feature-style explanations, with far less work on example-based explanations (Keane and Kenny, 2019). Because people naturally reason by analogy Sørmo et al. (2005); Schank et al. (2014); Kolodner (2014), explaining predictions to specific training data is intuitively appealing. Example-based explanations also connect to work on case-based reasoning Aamodt and Plaza (1994) by relating new problem instances to similar past ones, a problem-solving strategy people naturally use in decision-making (Newel and Simon, 1972). Rudin et al. (2022) thus argues for developing modern case-based methods as one of the grand challenges in interpretable machine learning.
In this work, we synthesize existing black-box and white-box methods toward building (training data) attributable-by-design models. Specifically, we introduce the wrapper box pipeline to combine the accuracy of modern neural models with the faithful, example-based explanations of classic models. Our approach effectively “wraps” a given neural model with one or more transparent classic models to maintain neural performance while improving interpretability. Building on the tradition of fitting fully connected layers on neural representations Krizhevsky et al. (2012); Simonyan and Zisserman (2015); He et al. (2016); Devlin et al. (2019), we fit white classic models on extracted neural representations for inference. The reasoning process of a classic wrapper model can then be faithfully explained by showing the specific training examples that led to each prediction (Figure 1).
Note that wrapper boxes do not attempt to approximate the underlying neural model, i.e., wrapper box vs. neural model predictions can differ. Our claim is that wrapper box predictions can be faithfully explained and that the predictive performance of wrapper boxes is largely comparable to the underlying neural models (see below).
@@
We also evaluate the effectiveness of wrapper boxes using representations from modern large language models (LLMs). Namely, we experiment with Llama 2-7B Instruct (Touvron et al., 2023), Llama 3-8B-Instruct (Dubey et al., 2024), Mistral-7B Instruct Jiang et al. (2023), and Gemma-7B Instruct Mesnard et al. (2024). Results show that wrapper boxes using zero-shot LLM representations strongly outperform baseline LLM performance across both tasks.
@@
Finally, we evaluate another use case: enabling model decisions to be contested based on the training data responsible for those decisions. Specifically, we consider identifying which training data needs to be removed to change a model decision Yang et al. (2023). This task offers a form of algorithmic recourse (Karimi et al., 2022), which emphasizes providing actionable explanations to users unfavorably treated by automated systems. Users are provided a foundation for contesting model decisions by attributing model decisions to specific training data. Compared to Yang et al. (2023), we show higher coverage and correctness in identifying which training data to remove while also generalizing beyond simple linear models and scaling to more modern neural networks.
@@ 2 Related Work
Case Influential Examples Explanation Examples Performance Faithfulness Simplicity
@@ III Subset Subset $\circ$ $\checkmark$ $\checkmark$
Table 2: Case-based models permit tradeoffs between key outcome variables – predictive performance, explanation faithfulness (or fidelity), and explanation simplicity – based on which (influential) training examples are used in making a prediction vs. to explain that prediction. Note that for any given input, different models will naturally vary in which training examples are influential in performing inference for that input.
@@ 3 Example-based Explanation Tradeoffs
We focus on interpretable predictive models that tie inference directly to specific training examples, enabling each prediction to be faithfully explained via those same training examples that determined the model’s prediction. Appendix D further discusses user perceptions of machine-retrieved examples. color=blue!20]Sam: appendix mention here
To better elucidate the design space for working with such models, this section illustrates possible tradeoffs between three key variables of interest: predictive performance, explanation faithfulness, and explanation simplicity. Following Jacovi and Goldberg (2020), we conceptually define faithfulness as how accurately presented explanations reflect the actual reasoning process of the inference model. Concretely, we evaluate the faithfulness of example-based explanations by completeness (Gu et al., 2023), where derived examples are faithful to the extent that all instances that support the test prediction are selected. The simplicity of example-based explanations can be intuitively quantified as the number of presented instances (Nguyen and Martínez, 2020).
Section 4 discusses how such tradeoffs can be operationalized in practice for our specific wrapper box models.
@@ 4 Wrapper Boxes
Our wrapper box pipeline essentially “wraps” a given neural model with one or more white box classic models fitted on extracted neural representations for inference. Note that the resultant classic models do not attempt to approximate the underlying neural model faithfully. While both classifiers leverage learned linearly separable neural representations, the underlying decision-making process differs. Hence, derived example-based explanations faithfully explain the inference procedure of the wrapper boxes (interpretable classic models), not the original neural model.
Post hoc methods evaluate fidelity for the neural model they seek to explain DeYoung et al. (2020); Jacovi and Goldberg (2020) since explanations can diverge from actual model behavior. In contrast, we leverage case-based classifiers where derived example-based explanations by construction must have been consulted during inference. Loss in fidelity can only occur intentionally if fewer training examples are shown in the explanation to reduce information overload.
@@ 4.1 Learning Feature Representations
As shown in Figure 1, we start with a fine-tuned neural model that acts as a task-specific encoder to learn high-quality embeddings for the input text. Whereas traditional neural models often fit linear classifiers on learned representations, we extract these representations for use by various classic, white box classifiers. This substitution thus enables prediction supported by faithful, example-based explanations and is agnostic to the neural architecture, training procedure, and data used.
@@ 4.2 Wrapper Box Models
We consider three case-based models in which inference is directly linked to training examples. This means that, by design, model predictions can be faithfully and intuitively attributed to specific relevant training examples.
@@ 5.1 Results
Results for large language models (LLMs)
Appendix H conducts an ablation study that evaluates the effectiveness of wrapper boxes using representations from modern LLMs. Namely, we experiment with Llama 2-7B-Instruct (Touvron et al., 2023), Llama 3-8B-Instruct (Dubey et al., 2024), Mistral-7B-Instruct Jiang et al. (2023), and Gemma-7B-Instruct Mesnard et al. (2024). Results show that wrapper boxes using zero-shot LLM representations strongly outperform baseline LLM performance across both tasks.
@@ 7 Conclusion
We propose wrapper boxes to provides faithful, example-based explanations for classic case-based model predictions, attributing decisions to specific training data. Our proposed pipeline is quite general and agnostic to the underlying neural architecture, training procedure, and input data. After training a neural model, the learned feature representation is input to white-box case-based reasoning models for prediction. Because case-based models tie inference directly to specific training data, each prediction can be faithfully attributed to the training examples responsible.
Our first evaluation showed that white case-based models could deliver predictive performance largely comparable to baseline transformers, as seen across seven large pre-trained language models, two datasets of varying scale, three classic models, and four metrics.
In addition, we discussed how such attribution enables automated decisions to be contested based on the training data responsible for those decisions. In comparison to prior work Yang et al. (2023), our approach achieves both higher coverage and correctness in identifying which training data to remove to change a model decision.
@@ Acknowledgments
This research was supported in part by Cisco and by Good Systems  [^1] http://goodsystems.utexas.edu/, a UT Austin Grand Challenge to develop responsible AI technologies. This work was also partially supported by NSF grant IIS-2107524. The statements made herein are solely the opinions of the authors and do not reflect the views of the sponsoring agencies.
@@ G.2 E-SNLI
Appendix H Wrapper Box Results for LLMs
H.1 Evaluation Setup
We additionally experiment with several more modern large language models (LLMs) to test the generalizability of wrapper boxes. Namely, we experiment with Llama 2-7B Instruct (Touvron et al., 2023), Llama 3-8B-Instruct (Dubey et al., 2024), Mistral-7B Instruct Jiang et al. (2023), and Gemma-7B Instruct Mesnard et al. (2024). We used model checkpoints publicly hosted on Hugging Face. Table 14 shows the prompts used for all LLMs.
Representations are extracted from the penultimate layer (directly preceding the language modeling heads) and mean-pooled across tokens to obtain a sentence-level embedding for white classic models. Sentence representations are then further processed by fitting a logistic regression, and we then take the logit output of the regression model to be the final input features for wrapper boxes. We empirically observe that this logistic transformation is necessary to achieve comparable performance and that fitting wrapper boxes directly on mean-pooled embeddings can lead to severe performance degradation, particularly for $L$-means.
We use the same datasets and metrics introduced in Section 5. Due to time and computation constraints, we only report zero-shot results and only report metrics on a 10,000 random stratified sample for E-SNLI. Applying state-of-the-art in-context-learning (ICL) strategies or fine-tuning may improve baseline neural performance but would also result in different representations as input to wrapper boxes. The efficacy of wrapper boxes in these scenarios would thus need to be empirically benchmarked, although we anticipate that there should be no drastic performance degradation.
H.2 ToxiGen LLM Results
Table 12 shows predictive performance results for ToxiGen. LLM zero-shot results favor recall over precision, likely due to the imbalanced distribution of labels in ToxiGen, with Llama 3 8B being the best model. No wrapper box uniformly outperforms others, although the precision and recall scores are more balanced. Results show that wrapper boxes using zero-shot LLM representations strongly outperform baseline LLM performance across both tasks.
Model Classifier Accuracy (%) F1 Score (%) Precision (%) Recall (%)
LLama2 7B Zero-shot 61.17 55.76 42.51 80.99
KNN 82.02 68.05 73.47 63.38
DT 80.85 69.90 66.56 73.59
L-Means 82.23 71.06 69.97 72.18
LLama3 8B Zero-shot 71.70 65.81 51.82 90.14
KNN 77.23 59.47 64.34 55.28
DT 76.70 61.10 61.65 60.56
L-Means 76.60 61.27 61.27 61.27
Mistral 7B Zero-shot 68.62 64.50 48.99 94.37
KNN 78.72 62.69 66.67 59.15
DT 79.04 63.59 66.93 60.56
L-Means 78.72 63.90 65.56 62.32
Gemma 7B Zero-shot 55.64 56.79 40.23 96.48
KNN 79.36 63.12 68.60 58.45
DT 78.94 68.37 62.57 75.35
L-Means 79.47 66.55 65.53 67.61
Table 12: ToxiGen test set performance metrics (percentage) of wrapper boxes for various instruction-tuned, open-source large language models. Embeddings are mean-pooled overall all sequence tokens from the penultimate layer (directly preceding the generative language modeling head). Then, representations are further transformed into logits using weights learned from logistic regressor before input as features to other wrapper boxes (KNN, DT, L-Means).
H.3 E-SNLI LLM Results
Table 13 shows predictive performance results for E-SNLI. Precision and recall scores are balanced for both LLMs and wrapper boxes here since E-SNLI has a balanced distribution of labels. Interestingly, we observe that decision tree consistently outperforms other wrapper boxes, although the differences (e.g., compared to KNN) may not be significant. Nevertheless, we consistently observe that wrapper boxes can strongly outperform baseline LLMs using zero-shot LLM representations.
Model Classifier Accuracy (%) F1 Score (%) Precision (%) Recall (%)
LLama2 7B Zero-shot 47.96 42.28 50.52 48.51
KNN 70.80 70.76 70.92 70.72
DT 73.03 72.91 72.91 72.96
L-Means 65.95 66.42 69.37 65.87
LLama3 8B Zero-shot 61.93 62.71 68.48 61.89
KNN 70.75 70.72 70.86 70.67
DT 73.35 73.31 73.34 73.30
L-Means 66.43 66.83 69.53 66.33
Mistral 7B Zero-shot 70.99 71.47 72.81 70.99
KNN 80.23 80.15 80.13 80.18
DT 81.94 81.90 81.91 81.91
L-Means 77.20 76.27 76.59 77.00
Gemma 7B Zero-shot 52.56 48.12 70.30 53.20
KNN 66.07 65.98 66.15 65.97
DT 69.68 69.68 69.80 69.62
L-Means 62.61 63.02 66.41 62.50
Table 13: E-SNLI test set performance metrics (percentage) of wrapper boxes for various instruction-tuned, open-source large language models. Embeddings are mean-pooled overall all sequence tokens from the penultimate layer (directly preceding the generative language modeling head). Then, representations are further transformed into logits using weights learned from logistic regressor before input as features to other wrapper boxes (e.g., KNN, DT, L-Means). The training set is a 10,000 random stratified sample of the original  500k set.
Dataset Prompt
ToxiGen You are a content moderator tasked with toxicity detection. After reading the provided text, categorize it as either "0" for benign or "1" for toxic. Only respond with "0" or "1". Do not include any additional text or explanations.
E-SNLI You are an informative AI assistant tasked with natural language inference. Given the following premise and hypothesis, classify their relationship as "0" for entailment, "1" for neutral, or "2" for contradiction. Entailment (0): The hypothesis logically follows from the premise. If the premise is true, the hypothesis must also be true. Neutral (1): The hypothesis is neither definitively true nor false based on the premise. The premise provides some information, but it is insufficient to confirm or deny the hypothesis. Contradiction (2): The hypothesis directly conflicts with the premise. If the premise is true, the hypothesis must be false. Only respond with "0", "1", or "2". Do not include any additional text or explanations.
Table 14: LLM Prompts for ToxiGen and E-SNLI.